# Conformal Prediction with Upper and Lower Bound Models

### v1

@@ Abstract
This paper studies a Conformal Prediction (CP) methodology for building prediction intervals in a regression setting, given only deterministic lower and upper bounds on the target variable. It proposes a new CP mechanism (CPUL) that goes beyond post-processing by adopting a model selection approach over multiple nested interval construction methods. Paradoxically, many well-established CP methods, including CPUL, may fail to provide adequate coverage in regions where the bounds are tight. To remedy this limitation, the paper proposes an optimal thresholding mechanism, OMLT, that adjusts CPUL intervals in tight regions with undercoverage. The combined CPUL-OMLT is validated on large-scale learning tasks where the goal is to bound the optimal value of a parametric optimization problem. The experimental results demonstrate substantial improvements over baseline methods across various datasets.
footnotetext: Corresponding author: Miao Li (mli746@gatech.edu).
@@ 1 Introduction
Quantifying the uncertainty of machine learning models is crucial for numerous applications, particularly in large-scale real-world scenarios where prediction sets, rather than point predictions, enable more flexible and informed decision making. Uncertainty quantification (UQ) methods are essential for characterizing the unpredictibility arising in various real-world problems across science and engineering. Initially proposed by Vovk et al. (2005), CP is a popular distribution-free method for UQ, largely due to its ability to provide finite-sample coverage guarantees and its computational efficiency.
Most studies in CP focus on constructing prediction intervals based on a fitted mean model. This work introduces a novel setting where the value of interest is estimated using only a pair of valid upper and lower bounds, instead of a mean model. While valid bounds themselves provide perfect coverage by definition, they can sometimes be overly conservative. By slightly reducing the coverage level, these bounds can be tightened, resulting in significantly smaller intervals with theoretical guarantees and greater utility for decision making. The task of identifying valid upper and lower bounds arises in numerous applications, including estimating and bounding the frequency of a queried object using sketches charikar2002finding, a foundational problem relevant to modern machine learning goyal2012sketch; cormode2018privacy; zhang2014these; evaluating best- and worst-case scenarios in policy optimization feng2021non, investment analysis kamma2022near; wu2023stability; ma2017dual, power consumption for resource allocation albert2013inference, and so on.
In real-life applications, these bounds can often be obtained by formulating optimization problems. Specifically, upper and lower bounds on the optimal value can be certified through primal- and dual-feasible solutions feng2021non; ma2017dual; cappart2019improving; bergman2014optimization; kamma2022near. By leveraging the duality theorem, a primal feasible solution provides a valid upper bound, while a dual feasible solution yields a valid lower bound for the true optimal value. An exciting recent research direction has explored how to learn valid bounds for parametric optimization problems, where machine learning models (e.g., primal and dual optimization proxies chen2024real; qiu2024dual; demelas2024predicting; tanneau2024dual) predict close-to-optimal solutions that are guaranteed to be feasible by construction.
However, valid bounds may sometimes be too loose to offer meaningful insights, prompting efforts to refine them. For instance, in the frequency estimation problem using sketches, cormode2005improved introduced a deterministic upper bound to derive confidence intervals, which has been further refined in subsequent studies. Notably, sesia2023conformal enhances the efficiency of these intervals using CP by incorporating a natural lower bound of zero alongside the deterministic upper bound, enabling more precise UQ for frequency estimation, effectively capturing the structure of the problem at hand.
Beyond the sole purpose of improving and providing more informative bounds, refined prediction intervals can play a pivotal role in algorithmic design and modeling. Recent studies highlight the integration of UQ into model development, demonstrating its potential to enhance algorithmic performance liang2023conformal; chee2024uncertainty; abdar2021review.
This paper introduces CPUL, a novel CP mechanism specifically designed for this upper and lower bound setting. By integrating multiple interval construction strategies within the nested CP framework, CPUL maximally exploits the structure of the bounds. In this setting, many established CP methods paradoxically struggle to maintain adequate coverage in regions where the bounds are most tight. To overcome this same limitation in CPUL, a powerful optimal threshold mechanism (OMLT) is introduced to strategically refine CPUL intervals. The combined CPUL-OMLT reliably produces state-of-the-art results on extensive experiments on large-scale learning tasks for bounding the optimal value of a parametric optimization problem arising in energy systems.
The contributions of the paper can be summarized as follows: (1) it studies a new setting for CP where only pair of valid upper and lower bound models, instead of a mean model, is available for the quantity of interest; (2) it proposes the CPUL mechanism to fully leverage the information provided by the given bounds, delivering efficient and reliable predictions; and (3) it augments CPUL with the OMLT mechanism to address the notably higher miscoverage rates in regions with initially tight bounds.
@@ 1.1 Problem Statement
Consider a random variable $X$ with support in $\mathbb{R}^{d}$, and a real Borel measurable function $f\,{:}\,\mathcal{X}\rightarrow\mathbb{R}$, which represents the unknown underlying true function that determines the label $Y=f(X)\in\mathbb{R}$. Let $\hat{B}^{l}$ and $\hat{B}^{u}$ be lower and upper predictors for $f$, i.e.,
@@
It immediately follows that $[\hat{B}^{l}(X),\hat{B}^{u}(X)]$ yields a prediction interval for $Y$ with 100% coverage. The goal of this paper is to derive tighter intervals for $Y$ by accepting a $1-\alpha$ marginal coverage guarantee for $\alpha\in[0,1]$.
Formally, consider a dataset of $N$ independent and identically distributed (i.i.d.) observations $(X_{i}\,{,}\,Y_{i})_{i\in[N]}$, where $X_{i}$ is sampled from an underlying distribution $\mathcal{P}$, and $Y_{i}\,{=}\,f(X_{i}),\forall i\,{\in}[N]$. Following the standard procedure of the split CP framework vovk2005algorithmic, divide the $N$ samples into a training and a calibration set, denoted as $\mathcal{D}_{\text{train}}$ and $\mathcal{D}_{\text{cal}},$ with corresponding indices $\mathcal{I}_{\text{train}},\mathcal{I}_{\text{cal}}$. Next, consider a new i.i.d. observation $X_{N+1}$, sampled from $\mathcal{P}$. For $\alpha\,{\in}\,[0,1]$, the goal is to refine the initial prediction interval
$\tilde{C}(X_{N+1})=\left[\hat{B}^{l}(X_{N+1}),\ \hat{B}^{u}(X_{N+1})\right]$ (2)
into a (smaller) prediction interval
$\hat{C}(X_{N+1})=\left[\hat{L}(X_{N+1}),\ \hat{U}(X_{N+1})\right]$ (3)
that satisfies the marginal coverage guarantee
@@
Indeed, recall that, although $\tilde{C}(X_{N+1})$ has 100% coverage by construction, it may be too wide to derive meaningful insights. This motivates the need to derive smaller prediction intervals while ensuring reasonable marginal coverage.
The above goal can be stated as the optimization problem
@@
where $\hat{C}$ maps $X\in\mathbb{R}^{d}$ to a subset of $\mathbb{R}$, and the length of a prediction interval $C$ is defined as
[…]
where $\mathbf{1}_{\mathcal{X}}$ is the indicator function of $\mathcal{X}$.
Remark. The setting naturally extends to the case where multiple lower bounds $\{\hat{B}^{l}_{1},\ldots,\hat{B}^{l}_{B_{l}}\}$ and upper bounds $\{\hat{B}^{u}_{1},\ldots,\hat{B}^{u}_{B_{u}}\}$ are provided. In this case, define
$\hat{B}^{l}(x)=\max\{\hat{B}^{l}_{1}(x),\ldots,\hat{B}^{l}_{B_{l}}(x)\},$
$\hat{B}^{u}(x)=\min\{\hat{B}^{u}_{1}(x),\ldots,\hat{B}^{u}_{B_{u}}(x)\},$
which yields valid lower and upper bound predictors that leverage all available information effectively.
@@ 2 Background
This section presents several standard CP techniques such as Split Conformal Prediction, Conformal Quantile Regression, and Nested Conformal Prediction. It is important to note that, although these techniques can be applied to the problem at hand (see experiment details in Section 5 and Appendix A.4.1), unlike CPUL, they are not designed to explicitly leverage valid lower and upper bound predictors.
Split Conformal Prediction (SCP) vovk2005algorithmic; papadopoulos2002inductive; lei2018distribution is one of the most commonly used CP frameworks. Therein, a mean prediction model $\hat{f}$ is first trained using the training data $\mathcal{D}_{train}$. SCP then produces prediction intervals of the form
$\tilde{C}_{\text{SCP}}(x)=\left\{y\,\middle|\,\tilde{Q}^{\text{SCP}}_{\frac{\alpha}{2}}\leq\hat{s}(x,y)\leq\tilde{Q}^{\text{SCP}}_{1-\frac{\alpha}{2}}\right\},$ (7)
where $\hat{s}\,{:}\,\mathbb{R}^{d}{\times}\mathbb{R}\,{\rightarrow}\,\mathbb{R}$ is a so-called conformity score, and $\tilde{Q}^{\text{SCP}}_{\frac{\alpha}{2}},\tilde{Q}^{\text{SCP}}_{1-\frac{\alpha}{2}}$, denote the empirical $\frac{\alpha}{2}$ and $1{-}\frac{\alpha}{2}$ quantiles of the scores $\{\hat{s}(X_{i},Y_{i})\}_{i\in\mathcal{I}_{\text{cal}}}$. Under the exchangeability assumption, the resulting prediction intervals have valid marginal coverage vovk2005algorithmic, i.e.,
@@
Conformal Quantile Regression (CQR) To alleviate SCP’s lack of local adaptivity, CQR romano2019conformalized; sesia2020comparison combines a quantile regression model with a conformalization procedure as follows. CQR requires that a quantile regression model is first trained to predict the quantiles $\alpha_{\text{lo}}$ and $\alpha_{\text{hi}}$. Then, conformalized prediction intervals are of the form
$\hat{C}^{\text{CQR}}(x)=\left[\hat{q}_{\alpha_{\text{lo}}}(x){-}\tilde{Q}^{\text{CQR}}_{1-\alpha},\ \hat{q}_{\alpha_{\text{hi}}}(x){+}\tilde{Q}^{\text{CQR}}_{1-\alpha}\right],$ (9)
where $\hat{q}_{\alpha_{\text{lo}}}(x),\hat{q}_{\alpha_{\text{hi}}}(x)$ are the predicted lower and upper quantiles given input data $x$, and $\tilde{Q}^{\text{CQR}}_{1-\alpha}$ is the $1{-}\alpha$ quantile of the conformity scores $\{\hat{s}^{\text{CQR}}(X_{i},Y_{i})\}_{i\in\mathcal{I}_{\text{cal}}}$, defined as
@@
The reader is referred to sesia2020comparison for variants of the original CQR construction.
The CQR method could be applied to the setting at hand by approximating $\hat{q}_{\alpha_{\text{lo}}},\hat{q}_{\alpha_{\text{hi}}}$ with $\hat{B}^{l},\hat{B}^{u}$ in the above derivation. However, experiments demonstrate that this approach does not adequately capture the heteroscedasticity of residuals between the lower- and upper-bound predictors. An efficient mechanism to alleviate this issue is presented in Section 3, which does not require training any additional model.
Nested Conformal Prediction (NCP) gupta2022nested introduces a unifying CP framework using nested prediction sets. The NCP framework considers a family of nested prediction intervals $\{\hat{C}_{t}\}_{t\in\mathcal{T}\subseteq\mathbb{R}}$, i.e.,
@@
Other Related Methods sesia2023conformal construct prediction intervals for the frequency of a queried object based on a sketch charikar2002finding. Therein, a deterministic upper bound $\hat{B}^{u}$ provided by cormode2005improved is combined with a trivial lower bound $\hat{B}^{l}=0$ (reflecting the non-negativity of counts) to construct nested intervals. The construction, referred to as “SFD CP” in this paper, is defined as:
@@ 3 CP With Upper And Lower Bounds
This section presents CPUL, which is specifically designed to exploit the strong properties of $\hat{B}^{l},\hat{B}^{u}$. Indeed, the main limitation of traditional CP methods such as SCP and CQR is that they do not leverage the fact that $\hat{B}^{l},\hat{B}^{u}$ yield valid lower and upper bounds on the target variable. Additionally, these methods do not necessarily integrate information from both models and are not inherently designed to account for the heteroskedasticity of residuals between these bounds (e.g., $\hat{B}^{l}$ might provide accurate estimates while $\hat{B}^{u}$ could perform poorly). These observations are further validated in the results of Section 5.

### v2

@@ Abstract
Decision makers routinely use optimization technology to plan and operate complex systems like global supply chains or power grids. In that context, practitioners face the challenge of assessing the optimality gap of a given solution, i.e., how optimal or sub-optimal said solution is. While constrained optimization algorithms generally provide such information in the form of so-called dual bounds, these bounds are often too loose to be informative. Therefore, decision makers must evaluate solution quality using their domain knowledge and experience, which is prone to errors and may lead to sub-optimal decisions. This paper addresses that challenge by providing estimates of solution quality in the form of prediction intervals, obtained via conformal prediction techniques. To that end, the paper introduces a general framework for conformal prediction with upper and lower bound predictors, which explicitly addresses the potential heteroskedasticity of either predictor via a model selection step. The paper also proposes a principled mechanism to avoid paradoxical miscoverage in cases where upper/lower predictors are both highly accurate. Finally, numerical experiments on large-scale problems found in industry demonstrate that the proposed approach achieves substantial reduction in interval width over several baseline methods.
@@ 1 Introduction
Optimization technology is instrumental in operating complex systems efficiently, from supply chains, logistics and manufacturing to healthcare and power systems. Nevertheless, it is often intractable to obtain optimal solutions to large-scale, nonlinear and non-convex problems found in industry. This limitation has motivated the use of approximate and heuristic algorithms, which are fast but may return sub-optimal solutions. As a consequence, decision makers face the hurdle of assessing whether a given solution is optimal enough, or whether substantial improvements are possible. This assessment requires computing dual bounds, i.e., mathematical certificates of solution quality, which are typically obtained by solving relaxations boyd2004convex; geoffrion2009lagrangean or by branch-and-bound algorithms wolsey1999integer. However, in real-life applications, these dual bounds may be too loose to be informative, thus leaving practitioners to evaluate solution quality by relying on their domain knowledge and experience, which is prone to errors and may lead to sub-optimal decisions.
To address this challenge, this paper proposes a data-driven approach that evaluates the sub-optimality of a solution by leveraging uncertainty quantification (UQ) and conformal prediction (CP) techniques. Namely, the paper proposes to provide probabilistic guarantees for solution quality in the form of prediction intervals. To the best of the authors’ knowledge, this paper is the first to propose the use of CP to estimate solution sub-optimality in an optimization setting. In doing so, the paper introduces CPUL, a general framework for CP with upper and lower bounds. This contrasts with existing approaches from the optimization literature, which almost exclusively rely on computing valid, but possibly loose, dual bounds. The paper’s setting also contrasts with traditional CP settings that rely on approximate predictors, e.g., mean predictors vovk2005algorithmic or quantile regression models romano2019conformalized which, in the present setting, may not yield valid lower/upper bounds.
The paper’s contributions are summarized as follows: (1) it presents a data-driven methodology to evaluate the optimality of solutions to optimization problems using uncertainty quantification techniques, (2) it introduces CPUL, a general framework for CP with upper and lower bounds, with a focus on capturing the heteroskedasticity of lower and upper bound residuals, (3) it addresses the paradox of miscoverage depsite high-quality lower and upper bounds by proposing an additional mechanism, Optimal Minimum Length Threshold (OMLT), (4) it conducts numerical experiments on large-scale economic dispatch problems used in real-time electricity markets.
@@ 1.1 Problem statement
Motivating Example
Consider a parametric optimization problem of the form
$\Phi(x)=\min_{s}\quad$ $f_{x}(s)$ (1a)
s.t. $g_{x}(s)\leq 0,$ (1b)
where $x\in\mathcal{X}\subseteq\mathbb{R}^{p}$ denotes the instance’s parameters, $s\in\mathbb{R}^{n}$ are decision variables, $f_{x}:\mathbb{R}^{m}\rightarrow\mathbb{R}$ is the objective function to be minimized, $g_{x}:\mathbb{R}^{n}\rightarrow\mathbb{R}^{m}$ encodes constraints and $\mathcal{S}_{x}=\{g_{x}(s)\leq 0\}$ is the set of feasible solutions, and $\Phi(x)$ is the instance’s optimal value. The paper assumes the availability of a primal-dual pair in the following form. A primal solution $\bar{s}\in\mathcal{S}_{x}$ provides a primal (upper) bound on the optimal value, i.e., $f_{x}(\bar{s})\geq\Phi(x)$. Primal solutions are obtained using either heuristics of exact algorithms. Conversely, a dual (lower) bound $\psi\in\mathbb{R}$ is guaranteed to be smaller than the optimal value, i.e., $\psi\leq\Phi(x)$. Dual bounds are obtained from relaxations and/or branch-and-bound algorithms, and provide a mathematical certificate on the quality of a primal solution $\bar{s}$. It is important to note that all modern global optimization algorithms return such a primal-dual pair. It is also possible to obtain a primal-dual pair using only machine learning (ML) models, by combining primal-feasible ML architectures donti2021dc3; chen2023end with the dual lagrangian learning framework introduced in tanneau2024dual.
General problem setting
Consider a random variable $X$ with support $\mathcal{X}\subseteq\mathbb{R}^{d}$, and an unknown Borel measurable function $f{:}\,\mathcal{X}\rightarrow\mathbb{R}$ that defines a target $Y\,{=}\,f(X)\,{\in}\,\mathbb{R}$. Let $\hat{B}^{l}$ and $\hat{B}^{u}$ be given lower and upper predictors for $f$, i.e.,
@@
These predictors inherently form a 100% prediction interval $[\hat{B}^{l}(X),\hat{B}^{u}(X)]$ for $Y$. Recall that, in a constrained optimization context, upper and lower predictors are obtained from primal and dual information, respectively. The goal of this paper is to derive tighter prediction intervals by targeting a $(1-\alpha)$ marginal coverage for $Y$ (where $\alpha\,{\in}\,[0,1]$), rather than perfect coverage.
Formally, consider a dataset of $N$ independent and identically distributed (i.i.d.) observations $(X_{i}\,{,}\,Y_{i})_{i\in[N]}$, where $X_{i}$ is sampled from an underlying distribution $\mathcal{P}$, and $Y_{i}\,{=}\,f(X_{i}),\forall i\,{\in}[N]$. Following the standard procedure of the split CP framework vovk2005algorithmic, divide the $N$ samples into a training and a calibration set, denoted as $\mathcal{D}_{\text{train}}$ and $\mathcal{D}_{\text{cal}},$ with corresponding indices $\mathcal{I}_{\text{train}},\mathcal{I}_{\text{cal}}$. Next, consider a new i.i.d. observation $X_{N+1}$, sampled from $\mathcal{P}$. For $\alpha\,{\in}\,[0,1]$, the goal is to refine the initial prediction interval $\tilde{C}(X_{N+1})=[\hat{B}^{l}(X_{N+1}),\hat{B}^{u}(X_{N+1})]$ into a (smaller) prediction interval $\hat{C}(X_{N+1})=[\hat{L}(X_{N+1}),\hat{U}(X_{N+1})]$ that satisfies the marginal coverage guarantee
@@
Indeed, recall that, while $\tilde{C}(X_{N+1})$ has 100% coverage by construction, it may be too wide to derive meaningful insights. This motivates the need to derive smaller prediction intervals while ensuring reasonable marginal coverage. Then the goal can be restated as the optimization problem
@@
where $\hat{C}$ maps $X\in\mathbb{R}^{d}$ to a subset of $\mathbb{R}$.
Remark 1. When multiple bounds $\{\hat{B}^{l}_{1},\ldots,\hat{B}^{l}_{B_{l}}\}$ and $\{\hat{B}^{u}_{1},\ldots,\hat{B}^{u}_{B_{u}}\}$ are provided, valid predictors $\hat{B}^{l}(x)$ and $\hat{B}^{u}(x)$ leveraging all this information effectively can be defined as
$\hat{B}^{l}(x)=\max\{\hat{B}^{l}_{1}(x),\ldots,\hat{B}^{l}_{B_{l}}(x)\},\quad\hat{B}^{u}(x)=\min\{\hat{B}^{u}_{1}(x),\ldots,\hat{B}^{u}_{B_{u}}(x)\}.$ (5)
@@ 2 Background
UQ is crucial for informed decision making, particularly in real-world applications where prediction sets are preferred over point estimates. CP, first proposed by vovk2005algorithmic, is a widely used distribution-free UQ method, valued for its finite-sample coverage guarantees and computational efficiency. This section presents standard CP techniques, including Split CP, Conformal Quantile Regression, and Nested CP. While these methods can be applied to the problem at hand (see Section 5 and Appendix A.4.1), unlike the proposed framework, they are not designed to explicitly leverage valid lower and upper bound predictors.
Split Conformal Prediction (SCP)
vovk2005algorithmic; papadopoulos2002inductive; lei2018distribution is one of the most commonly used CP frameworks. Therein, a mean prediction model $\hat{f}$ is first trained using the training data $\mathcal{D}_{train}$. SCP then produces prediction intervals of the form
$\hat{C}_{\text{SCP}}(x)=\left\{y\,\middle|\,\hat{Q}^{\text{SCP}}_{\frac{\alpha}{2}}\leq\hat{s}(x,y)\leq\hat{Q}^{\text{SCP}}_{1-\frac{\alpha}{2}}\right\},$ (6)
where $\hat{s}\,{:}\,\mathbb{R}^{d}{\times}\mathbb{R}\,{\rightarrow}\,\mathbb{R}$ is a so-called conformity score, and $\hat{Q}^{\text{SCP}}_{\frac{\alpha}{2}},\hat{Q}^{\text{SCP}}_{1-\frac{\alpha}{2}}$, denote the empirical $\frac{\alpha}{2}$ and $1{-}\frac{\alpha}{2}$ quantiles of the scores $\{\hat{s}(X_{i},Y_{i})\}_{i\in\mathcal{I}_{\text{cal}}}$. Under the exchangeability assumption, the resulting prediction intervals have valid marginal coverage vovk2005algorithmic, i.e.,
@@
Conformal Quantile Regression (CQR)
To alleviate SCP’s lack of local adaptivity, CQR romano2019conformalized combines quantile regression models with a conformalization procedure: First, CQR trains quantile regression models to predict the quantiles $\alpha_{\text{lo}}$ and $\alpha_{\text{hi}}$. Then, conformalize as follows
$\hat{C}^{\text{CQR}}(x)=\left[\hat{q}_{\alpha_{\text{lo}}}(x){-}\hat{Q}^{\text{CQR}}_{1-\alpha},\ \hat{q}_{\alpha_{\text{hi}}}(x){+}\hat{Q}^{\text{CQR}}_{1-\alpha}\right],$ (8)
where $\hat{q}_{\alpha_{\text{lo}}}(x),\hat{q}_{\alpha_{\text{hi}}}(x)$ are the predicted lower and upper quantiles given input data $x$, and $\hat{Q}^{\text{CQR}}_{1-\alpha}$ is the $1{-}\alpha$ quantile of the conformity scores $\{\hat{s}^{\text{CQR}}(X_{i},Y_{i})\}_{i\in\mathcal{I}_{\text{cal}}}$, defined as
@@
The CQR method could be adapted to the setting at hand by approximating $\hat{q}_{\alpha_{\text{lo}}},\hat{q}_{\alpha_{\text{hi}}}$ with $\hat{B}^{l},\hat{B}^{u}$ in the above derivation. Unlike the standard CQR, which requires fitting auxiliary quantile regression models whose accuracy dictates performance (and may require large training datasets in high-dimensional settings), $\hat{B}^{l},\hat{B}^{u}$ from the constrained optimization formulation offer key advantages. These bounds are often directly computable from the base model (e.g., once a feasible primal model is constructed, an upper bound can be computed by evaluating the objective at that solution), making them available at minimal cost. They also inherently guarantee perfect coverage, regardless of training data size. This paper focuses on effectively leveraging these readily available and valid bounds.
Nested Conformal Prediction (NCP)
gupta2022nested introduces a unifying CP framework by nested prediction sets. NCP considers a family of nested prediction intervals $\{\hat{C}_{t}\}_{t\in\mathcal{T}\subseteq\mathbb{R}}$, i.e.,
@@
Other Related Methods
CP techniques have been use to analyze the precision of sketching algorithms, which are increasingly vital tools for handling massive datasets in modern machine learning broder2004network; goyal2012sketch; cormode2018privacy; zhang2014these. sesia2023conformal construct prediction intervals for the frequency of a queried object based on a sketch charikar2002finding. Therein, a deterministic upper bound $\hat{B}^{u}$ provided by cormode2005improved is combined with a trivial lower bound $\hat{B}^{l}=0$ (reflecting the non-negativity of counts) to construct nested intervals. The construction, referred to as “SFD CP” in this paper, is defined as:
@@ 3 CP With Upper And Lower Bounds
(a) Illustration of the construction of prediction intervals in CPUL: (a) $\hat{C}^{\text{ll}}_{0}$, (b) $\hat{C}^{\text{uu}}_{0}$, (c) $\hat{C}^{\text{lu}}_{0}$, (d) $\hat{C}^{\text{ul}}_{0}$. Each family of prediction intervals is conformalized following the NCP framework (see Eq. (11)).
(b) Illustration of Paradoxical Miscoverage (motivation for OMLT): using a constant offset (±$\pm t$) in the NCP construction results in an empty prediction interval where $\hat{B}^{u}(x)-\hat{B}^{l}(x)$ is small.
Figure 1: Illustration of CPUL-OMLT construction
This section presents CPUL, designed to exploit the strong properties of $\hat{B}^{l},\hat{B}^{u}$. A key limitation of traditional CP methods like SCP and CQR is not leveraging that $\hat{B}^{l},\hat{B}^{u}$ yield valid lower and upper bounds on the target variable. Additionally, these methods do not necessarily integrate information from both models nor are they inherently designed to account for heteroskedastic residuals between these bounds (e.g., $\hat{B}^{l}$ might provide accurate estimates while $\hat{B}^{u}$ performs poorly). The results in Section 5 further validate these observations.